# Dynamic Inertial Poser (DynaIP): Part-Based Motion Dynamics Learning for Enhanced Human Pose Estimation with Sparse Inertial Sensors

### 02 Dec 2023 v1

@@ Abstract
This paper introduces a novel human pose estimation approach using sparse inertial sensors, addressing the shortcomings of previous methods reliant on synthetic data. It leverages a diverse array of real inertial motion capture data from different skeleton formats to improve motion diversity and model generalization. This method features two innovative components: a pseudo-velocity regression model for dynamic motion capture with inertial sensors, and a part-based model dividing the body and sensor data into three regions, each focusing on their unique characteristics. The approach demonstrates superior performance over state-of-the-art models across five public datasets, notably reducing pose error by 19% on the DIP-IMU dataset, thus representing a significant improvement in inertial sensor-based human pose estimation. We will make the implementation of our model available for public use.
@@ A Implementation Details
Training Details We train our model on the Pytorch platform, optimized by the Adam optimizer. The training data is segmented into chunks, each consisting of 300 frames. We conduct the training with a batch size of 256 and an initial learning rate of 0.001, which is adjusted using a cosine annealing scheduler. In the training phase, we perform training for 200 epochs with a GTX 3090 GPU, which takes around 2.5 hours.

### 07 Mar 2024 v2

@@ Abstract
This paper introduces a novel human pose estimation approach using sparse inertial sensors, addressing the shortcomings of previous methods reliant on synthetic data. It leverages a diverse array of real inertial motion capture data from different skeleton formats to improve motion diversity and model generalization. This method features two innovative components: a pseudo-velocity regression model for dynamic motion capture with inertial sensors, and a part-based model dividing the body and sensor data into three regions, each focusing on their unique characteristics. The approach demonstrates superior performance over state-of-the-art models across five public datasets, notably reducing pose error by 19% on the DIP-IMU dataset, thus representing a significant improvement in inertial sensor-based human pose estimation. Our codes are available at https://github.com/dx118/dynaip
@@ C Additional Analysis
Additional Comparison on Model performance. To isolate data impact and assess model performance, we provide a additional comparison with our model marked as ${}^{\dagger}$ using a data set consistent with PIP (combined with AMASS and DIP data) in Tab. 8. The performance enhancement highlights our architecture’s efficiency.
Dataset Method SIP Err(°) Ang Err(°) Pos Err(cm) Mesh Err(cm)
DIP_IMU PIP 15.02 8.73 5.04 5.95
${}^{\dagger}$ 14.11 7.00 4.97 5.97
Total Capture PIP 12.93 12.04 5.61 6.51
${}^{\dagger}$ 12.42 11.06 5.11 5.79
Table 8: The performance comparison on DIP-IMU and Totalcapture. ${}^{\dagger}$ ia trained on AMASS and DIP-IMU, same as PIP.
(a) DynaIP*
(b) DynaIP*-Position
(c) PIP-official
Figure 11: Visualization of first linear layer weights, Accs on marked cols while the rest are Oris. (PIP Accs on col 0 to 17)
@@ D Additional Qualitative Results
Effect of using velocity as intermediate outputs. Previous works, such as DIP, identified a trend where networks tend to discard much of the acceleration, and introduced an auxiliary task to reconstruct the acceleration for alleviating this problem. To better utilize acceleration, we first regress the pseudo velocity as an intermediary output, supervised by the ground-truth velocity obtained from the human pose. Compared with PIP/Transpose that uses leaf joints positions as intermediate outputs, our method has certain advantages in terms of using acceleration. Since noise in acceleration easily diverges with quadratic integration, it’s more difficult to capture the spatial relationship in noisy acceleration. Instead, the network learns to infer the position more from the orientation input due to human kinematic relationship. Following DIP’s explanation, we visualize the weights of the first linear layer of the network in the Fig. 11, where both the weights of PIP and ours-w.position in the corresponding dimension of the acceleration input is almost zero.